# DeepPAAC: A New Deep Galerkin Method for Principal-Agent Problems

### v1

@@ 4.5 Multidimensional Problem with Multidimensional Controls
The Hamiltonian is quadratic in each of $a$ and $Z$ however note the high-order term $0.5a^{2}Z^{2}V^{P}_{ww}$. The first order conditions for each control are:
${\cal F}_{Z}(t,x,w):$ $=a^{2}zV^{P}_{ww}+a^{2}V^{P}_{xw};$
${\cal F}_{a}(t,x,w):$ $=V^{P}_{x}+a(V^{P}_{xx}+Z^{2}V^{P}_{ww})+2aZV^{P}_{xw}.$
@@
In Figure 9 (a)-(d), it is seen that the Principal’s value function $V^{P}$ increases from upper left to lower right, meaning that it is increasing in $x$ and decreasing in $w$. This is consistent with the theoretical result that the Principal’s value function increases when the initial value of output process $\{X_{t}\}$ increases or the Agent’s participant constraint $R$ decreases. Figure 9(b)-(e) show that $Z={\partial V^{A}\over\partial x}$ decreases in both $x$ and $w$, which is consistent with the concavity condition for the Agent’s value function: ${\partial^{2}V^{A}\over\partial x^{2}}<0$. The latter also implies that ${\partial^{2}V^{A}\over\partial xw}<0$ for this example. Figure 9 (c) shows that the Agent’s optimal effort process $a^{*}$ (or equivalently, $\partial^{2}V^{A}\over\partial x^{2}$}) has a peak in the middle of the state space, maximized at moderate positive $x\simeq 0.5$ and moderate negative $w\simeq-1.5$. This peak shifts to a lower $w$ when we change to $\gamma_{P}=2$ (panel f).
@@ 5 Conclusion
and the Agent’s optimal effort process is $a^{*}_{t}=Z_{t},\forall 0\leq t\leq T$. When the Agent’s participation constraint is $R_{t}$ at time $t$, we have $W_{t}=-{1\over\gamma_{A}}\log(-R_{t})$.
@@
When Agent’s participant constraint is $R_{t}$ at time $t$, we have $x={1\over\gamma_{A}}\log(-R_{t})$. This yields the HJB equation for the Principal:
@@
$V^{P}(t,x)=\Phi^{P}\left(x+[{1\over 2}{(1+\gamma_{P})^{2}\over 1+\gamma_{A}+\gamma_{P}}-{\gamma_{P}\over 2}](T-t)\right)$

### v2

@@ 4.5 Multidimensional Problem with Multidimensional Controls
The Hamiltonian is quadratic in each of $a$ and $Z$ however note the high-order term $0.5a^{2}Z^{2}V^{P}_{ww}$. This yields the 2-dim. convergence loss criterion (taking the gradient in $a,Z$ respectively):
$L_{ctrl}=\Big[V^{P}_{x}+a(V^{P}_{xx}+Z^{2}V^{P}_{ww})+2aZV^{P}_{xw},\;a^{2}zV^{P}_{ww}+a^{2}V^{P}_{xw}\Big].$
@@
In Figure 9 (a)-(d), it is seen that the Principal’s value function $V^{P}$ increases from upper left to lower right, meaning that it is increasing in $x$ and decreasing in $w$. This is consistent with the theoretical result that the Principal’s value function increases when the initial value of output process $\{X_{t}\}$ increases or the Agent’s participant constraint $R$ decreases. Figure 9(b)-(e) show that $Z={\partial V^{A}\over\partial x}$ decreases in both $x$ and $w$, which is consistent with the concavity condition for the Agent’s value function: ${\partial^{2}V^{A}\over\partial x^{2}}<0$. The latter also implies that ${\partial^{2}V^{A}\over\partial x\partial w}<0$ for this example. Figure 9 (c) shows that the Agent’s optimal effort process $a^{*}$ (or equivalently, $\partial^{2}V^{A}\over\partial x^{2}$}) has a peak in the middle of the state space, maximized at moderate positive $x\simeq 0.5$ and moderate negative $w\simeq-1.5$. This peak shifts to a lower $w$ when we change to $\gamma_{P}=2$ (panel f).
@@ 5 Conclusion
and the Agent’s optimal effort process is $a^{*}_{t}=Z_{t},\forall 0\leq t\leq T$. When the Agent’s participation constraint is $R_{t}$ at time $t$, we have the initial condition $W_{t}=-{1\over\gamma_{A}}\log(-R_{t})$.
@@
When Agent’s participant constraint is $R_{t}$ at time $t$, the initial value is $x={1\over\gamma_{A}}\log(-R_{t})$. This yields the HJB equation for the Principal:
@@
$V^{P}(t,x)=\Phi^{P}\left(x+\Bigl[{1\over 2}{(1+\gamma_{P})^{2}\over 1+\gamma_{A}+\gamma_{P}}-{\gamma_{P}\over 2}\Bigr](T-t)\right)$